# Nonstabilizerness of a Boundary Time Crystal

### v1

@@ IV Nonstabilizerness in the mean-field limit
In the thermodynamic limit, the superradiant model of Eq. (2) behaves like a classical system: When $N\to\infty$, all correlations between collective spin variables become negligible and the system is exactly described by mean-field theory. This means that, in this limit, for the purpose of evaluating the expectation values of system observables, the system state can be approximated as a product state of single-particle density matrices,
@@
The key point is that, since correlations in collective observables are negligible in the thermodynamic limit, one is allowed to express the density matrix as a product state of individual components, when one deals with evaluation of expectations of observables. We will see in the following that this assumption may lead to incorrect predictions for the magic at finite sizes in some regions of the phase diagram, where the steady state is actually mixed. Indeed, the mean-field approximation is not well defined for mixed states, and this will have important consequences on the evaluation of quantum complexity measures.
@@
Using Eq. (6) and the fact that the state is a pure product state for $N\to\infty$, it is easy to compute the magic density of the average state $\hat{\rho}$ during its dynamics. In fact, we have $\mathcal{M}_{2}(\hat{\rho})=N\mathcal{M}_{2}(\hat{\rho}_{k})$, thus $\mathcal{M}_{2}(\hat{\rho})/N=m_{2}(\hat{\rho})=\mathcal{M}_{2}(\hat{\rho}_{k})$ and the latter is given by
@@
where the magnetization components obey Eqs. (14). We note that, in the thermodynamic limit, the density of magic is always finite, thus the stabilizer entropy always scales extensively with $N$, except when $m_{x}^{4}+m_{y}^{4}+m_{z}^{4}=1$. This only occurs when one of the magnetization components is equal to ±$\pm 1$ while the rest are zero, i. e., when the state is a stabilizer and is fully polarized along one of the three coordinate axes. We remark that Eq. (17) is only valid because correlations are negligible.
Figure 2: Density of (mean-field) magic as a function of time, obtained from Eqs. (14) and (17). (a) $\Omega=\omega_{0}/\kappa=0.5$. (b) $\Omega=1$. (c) $\Omega=2$. The initial condition is $\bm{m}(0)=(0,0,1)$.
Examples of magic dynamics in the thermodynamic limit are shown in Fig. 2 for different choices of the system parameters. The spin system is initialized in the fully polarized state $\bm{m}(0)=(0,0,1)$. We numerically integrate Eqs. (14) using implicit backward differentiation formulas of order 5 (BDF). Comparing the panels, we see that when $\Omega\leq\Omega_{c}$ the magic density quickly saturates to its stationary value after a transient time that depends on the initial state; on the other hand, for $\Omega>\Omega_{c}$, it displays never-ending oscillations akin to the magnetization oscillations in the BTC phase.
Interestingly, the magic density of the fixed point can be computed analytically using Eqs. (15) and (16). It reads
@@
It is shown in Fig. 3, blue line. At the critical point $\Omega=\Omega_{c}=1$, the fixed-point magic density is zero. This is because the fixed point in that case is the fully polarized eigenstate of $\hat{S}_{y}$, a stabilizer state. Moreover, the derivative of the magic density is singular at $\Omega=\Omega_{c}$, as $m_{2}$ has a cusp there, and has a finite jump $\Delta\overline{m_{2}}^{\prime}=4$. Thus, even if magic does not display any change of behavior between the magnetized and the time-crystal phase, it still shows singular behavior at criticality.
Figure 3: Density of magic of the fixed point [$\overline{m_{2}}$, Eq. (18)] and average density of magic along the closed orbits around the fixed point [$\langle\widetilde{m_{2}}\rangle$, Eq. (20)] in the mean-field limit ($\Omega=\omega_{0}/\kappa$). The black dotted line is the numerical fit of Eq. (22). The red dashed line indicates the saturation value, estimated using Eq. (IV).
The magic density is zero also for $\Omega=0$, where the fixed point is a steady state and is the dark state $|\ket{\psi_{D}}=\ket{\downarrow\downarrow\cdots\downarrow}$ (also a stabilizer state), and $\Omega\to\infty$, where fixed points are the eigenstates of $\hat{H}=\omega_{0}\hat{S}_{x}$. The maximum magic density is found for $\Omega^{*}_{1}=1/\sqrt{2}$ and $\Omega^{*}_{2}=\sqrt{2}$ and is equal to $\overline{m_{2}}^{*}=2\log 2-\log 3$, which coincides with the magic of the $T$ state |$\ket{T}\sim\ket{\uparrow}+\exp(i\pi/4)\ket{\downarrow}$.
However, it is important to note that, in the BTC phase, the fixed point of the classical MF dynamics is not a steady state, but an oscillation center in phase space [8]. Thus, in this phase it makes more sense to look at the average magic along the closed orbits encircling the fixed point, rather than the fixed point itself.
@@
On the other hand, this is not necessarily true for for magic. In this case, the average over the orbit of the limit cycle is not equivalent to the magic of the fixed point and is generally dependent on the initial condition, which determines the orbit in the semiclassical phase space.
To evaluate $m_{2}$ along these orbits, we proceed numerically. First, we generate $N_{\text{avg}}=1000$ initial conditions of the form $\bm{m}(0)=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, where $\theta\in[0,\pi]$ and $\phi\in[0,2\pi]$ are random variables uniformly distributed in their domains. For each of them, denoted $k$, we solve Eqs. (14) up to a long time $\tau$ and compute the instantaneous magic density $m_{2}^{(k)}(t)$ of this classical trajectory using Eq. (18). In our simulation, we consider $\kappa\tau=1000$. Then, we consider the (long-time) average magic along the trajectory,
@@
The result of this procedure is shown in Fig. 3 with orange scatter points. Error bars (denoting the standard error, i. e., the standard deviation divided by $\sqrt{N_{\text{avg}}}$) are not visible as they are smaller than the point size. We see that, in the magnetized phase $\Omega<1$, the magic computed by averaging over random initial states coincides with the magic of the stationary state, whereas for $\Omega>1$, in the BTC phase, the two quantities differ from each other. In either case, the magic density is finite, so the system always has extensive magic and there are no magic phase transitions associated with the mean-field transition. Yet, the cusp at criticality remains and even becomes more pronounced, as the right derivative of the averaged magic density at $\Omega_{c}$ becomes larger than the corresponding derivative of the fixed-point magic. Instead, the maximum at $\Omega_{2}^{*}=\sqrt{2}$ is gone, leaving the room to a monotonic growth of $\langle\widetilde{m_{2}}\rangle$ with the driving frequency.
A theoretical argument allows to compute its saturation value. When $\Omega\to\infty$, the oscillation frequency of the magnetization components goes to infinity as well. Thus, the three components will uniformly sample their respective domain and an average over time and initial conditions will be equivalent to an average over the solid angle of the magic density of the spherical vector $\bm{m}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$. Thus, one can estimate
@@ V Finite-size results
Figure 4: Density of magic of the average state evolving according to the Lindblad equation [Eq. (2)] as function of time, for several system sizes $N$, compared with the mean-field result. (a) $\Omega=\omega_{0}/\kappa=0.5$. (b) $\Omega=1$. (c) $\Omega=2$. The initial condition is the fully polarized state $|\ket{\psi(0)}=\ket{\uparrow\uparrow\dots\uparrow}$.
@@ V.1 Nonstabilizerness of the average state
We now turn our attention to the study of nonstabilizerness in finite-size systems. We start by directly solving the unconditioned Lindblad dynamics of Eq. (2), and we compute $\mathcal{M}_{2}$ of the average state according to Eq. (6). We integrate the Lindblad equation using the 4th order Runge-Kutta method. Due to the strong symmetry of the model, we can represent the entire dynamics in the maximum spin subspace ($S=N/2$), so the density matrix has a linear dimension $D=N+1$.
Our results are shown in Fig. 4, for system sizes up to $N=80$, for the same values of $\Omega$ shown in Fig. 2. In all panels, we see that the density of magic of the average state depends on the system size. For $\Omega\leq 1$, we see that the long-time density of magic of the average state coincides with the density of magic calculated using the mean-field approximation (the gray line), whereas for $\Omega>1$ this is not the case. In fact, in this regime the density of magic decreases with $N$ at long times, in contrast with the finite $m_{2}$ observed in MF dynamics.
@@
Let us consider the left-hand side of Eq. (23). If one takes the thermodynamic limit first, then the dynamics of collective observables is described by Eqs. (14), the system state is always pure, and the behavior of magic is the one shown in Fig. 2.
@@
The emergence of the time-crystal phase in the weakly dissipative regime is signaled by the fact that the amplitude of the magnetization oscillations decays more slowly as $N$ increases, but they are only persistent for $N\to\infty$: For any finite size, they will eventually decay, yielding a mixed steady state with $\lvert\bm{m}\rvert<1$ and a density of magic that goes to zero as $N$ grows.
Both these facts are not consistent with the mean-field limit, where $\hat{\rho}$ is a pure state with a finite density of magic. In particular, see Fig. 5, where we plot the purity of the average state for the same parameters as Fig. 4. Here is evident that, for any finite size, the purity of the average state is smaller than one in the BTC phase, as a consequence of the noncommutativity of the thermodynamic limit and the long-time limit, Eq. (23), and of the existence of residual correlations in the state.
Instead, the two limits in Eq. (23) do commute in the magnetized phase. Here, the purity goes to one, and the steady-state magic converges to its MF limit.
The behavior of the density of magic of the steady state is shown in Fig. 6(a) as a function of the driving frequency, further confirming that the Lindblad equation does not agree with the MF prediction of a finite density of magic for large $N$. This mismatch is accompanied by a sudden decrease in the purity of the steady state, as shown in Fig. 6(b), confirming the important role of correlations in this region for finite sizes. As long as the steady state is pure ($\Omega\leq 1$), the MF predictions and the magic of the finite-size state evolving through the Lindblad equation [Eq. (2)] agree with each other (compare this with the magic of the fixed point shown in Fig. 3 for $\Omega\leq 1$). This is no longer the case when $\text{Tr}\,\hat{\rho}^{2}<1$.
Figure 6: (a) Finite-size density of magic of the steady state for different system sizes, in a deterministic Lindblad evolution, as a function of $\Omega=\omega_{0}/\kappa$. (b) Purity of the steady state as a function of $\Omega$.
@@ V.2 Nonstabilizerness of unravelings
Figure 7: Finite-size long-time averaged magic density for different system sizes in different unravelings, see Eq. (5). (a) Quantum Jump ($\mu=0$). (b) $\mu=2$. (b) Quantum State Diffusion ($\mu\to\infty$). We also show the average mean-field magic $\langle\widetilde{m_{2}}\rangle$, Eq. (20) (orange points).
In this section, we show that the inconsistencies pointed out in the previous section disappear if, instead of computing the magic of the average state, one considers the average magic over ensembles of monitored quantum trajectories.
Indeed, in Fig. 7 we report the average long-time magic of ensembles of $M=500$ quantum trajectories, in three unravelings. In the three panels, orange scatter points represent the mean-field results discussed in Sec. IV. For $\Omega<1$ they correspond to the steady state, while for $\Omega>1$ to the average over time and random initial conditions. First of all, we see that the magic density in the $\Omega<1$ phase is essentially unchanged compared to Fig. 6, meaning that each quantum trajectory of the ensemble behaves very closely to the average state in this phase. This is also consistent with the early findings, showing that the variances of collective observables over the ensemble in this phase go to zero for large $N$ [52].
The situation is much different from the average one in the $\Omega>1$ phase, where we observe a finite, size-independent, density of magic that tells us that the magic of the ensemble trajectories is extensive, as opposed to the magic of the average state $\hat{\rho}$. Moreover, we see that the ensemble average of $m_{2}$, where each trajectory starts from the same initial state but follows a stochastic dynamics due to the random measurements performed by the environment, is essentially identical to the mean-field magic averaged over many initial states, Eq. (20) (see the orange points in Fig. 7). This observation tells us that the intrinsic randomness of quantum trajectories allows the system to explore its phase space, and that this information is completely hidden at the level of the average state evolving according to the deterministic Lindblad equation.
Moreover, we observe that the magic in all the unravelings is exactly the same, when the system size is sufficiently large, as we have analytically shown in Sec. III.1, in contrast with the behavior of the bipartite entanglement entropy [52]. We see that all the unravelings give the same magic except for finite-size corrections at small sizes, as shown in Fig. 7(b–c) for $\mu=2$ and $\mu\to\infty$.
@@
All these numerical results are consistent with the interpretation discussed in Sec. III.2: for large $N$, the SRE is unraveling-independent because correlations become negligible when computing the expectation values of collective observables. Indeed we can see by comparing the different panels of Fig. 7 that for small $N$ the magic density is not unraveling independent, due to correlations, and that unraveling-independence only kicks in for $N\gtrsim 20$ (empirically).
Thus, unraveling the dissipative dynamics in terms of quantum trajectories is able to resolve the inconsistency between finite-size simulations and the thermodynamic limit in the BTC phase, for two reasons. First of all, the states of the ensemble of quantum trajectories are all pure by construction, therefore the SRE is a proper complexity measure for each of them and there are no issues associated with nonpure states like for the magic of the Lindblad steady state in the BTC phase. Second, due to the intrinsic randomness of quantum trajectories, the dynamics of the single trajectories is much different from the dynamics of the Lindblad average state, and averaging over trajectories and time one gets a result well described in the large-size limit by the average over time and random initial states of MF trajectories.
@@
The same considerations apply to the density of magic as well, as shown in Fig. 8(b) for the same parameters. Also in this case, the magic dynamics along single trajectories more closely resembles the dynamics in the MF limit, compared to the Lindblad magic where instead oscillations decay over time and magic goes to zero. Moreover, since the SRE is a nonlinear function, its ensemble average over quantum trajectories does not coincide with the SRE of the average state. Therefore we think it is reasonable that its value, long-time averaged and also averaged over the pure states of the ensemble, would resemble the MF limit (averaged over many initial states) more than the magic of the average state for finite $N$.

### v3

@@ IV Nonstabilizerness in the mean-field limit
In the thermodynamic limit, the collective magnetizations of the superradiant model of Eq. (2) behave like those of a classical system: The correlations between collective spin variables scale more slowly with the system size $N$ than their expectations. So, when $N\to\infty$, the correlations become negligible compared to the expectations and the latter ones are exactly described by mean-field theory. This means that, in this limit, for the purpose of evaluating the expectation values of system observables, the system state can be approximated as a product state of single-particle density matrices,
@@
The key point is that, since correlations in collective observables are negligible in the thermodynamic limit, one is allowed to express the density matrix as a product state of individual components, when one deals with evaluation of expectations of observables. We will see in the following that this assumption may lead to incorrect predictions for the nonstabilizerness at finite sizes in some regions of the phase diagram, where the steady state is actually mixed. Indeed, the mean-field approximation is not well defined for mixed states, and this will have important consequences on the evaluation of quantum complexity measures.
@@
Using Eq. (6) and the fact that the state is a pure product state for $N\to\infty$, it is easy to compute the density of SRE of the average state $\hat{\rho}$ during its dynamics. In fact, we have $\mathcal{M}_{2}(\hat{\rho})=N\mathcal{M}_{2}(\hat{\rho}_{k})$, thus $\mathcal{M}_{2}(\hat{\rho})/N=m_{2}(\hat{\rho})=\mathcal{M}_{2}(\hat{\rho}_{k})$ and the latter is given by
@@
where the magnetization components obey Eqs. (14). We note that, in the thermodynamic limit, the density of SRE is always finite, thus the stabilizer entropy always scales extensively with $N$, except when $m_{x}^{4}+m_{y}^{4}+m_{z}^{4}=1$. This only occurs when one of the magnetization components is equal to ±$\pm 1$ while the rest are zero, i. e., when the state is a stabilizer and is fully polarized along one of the three coordinate axes. We remark that Eq. (17) is only valid because correlations are negligible.
Figure 2: Density of (mean-field) nonstabilizerness as a function of time, obtained from Eqs. (14) and (17). (a) $\Omega=\omega_{0}/\kappa=0.5$. (b) $\Omega=1$. (c) $\Omega=2$. The initial condition is $\bm{m}(0)=(0,0,1)$.
Examples of SRE dynamics in the thermodynamic limit are shown in Fig. 2 for different choices of the system parameters. The spin system is initialized in the fully polarized state $\bm{m}(0)=(0,0,1)$. We numerically integrate Eqs. (14) using implicit backward differentiation formulas of order 5 (BDF). Comparing the panels, we see that when $\Omega\leq\Omega_{c}$ the SRE density quickly saturates to its stationary value after a transient time that depends on the initial state; on the other hand, for $\Omega>\Omega_{c}$, it displays never-ending oscillations akin to the magnetization oscillations in the BTC phase.
Interestingly, the density of SRE of the fixed point can be computed analytically using Eqs. (15) and (16). It reads
@@
It is shown in Fig. 3, blue line. At the critical point $\Omega=\Omega_{c}=1$, the fixed-point nonstabilizerness density is zero. This is because the fixed point in that case is the fully polarized eigenstate of $\hat{S}_{y}$, a stabilizer state. Moreover, the derivative of the nonstabilizerness density is singular at $\Omega=\Omega_{c}$, as $m_{2}$ has a cusp there, and has a finite jump $\Delta\overline{m_{2}}^{\prime}=4$. Thus, even if nonstabilizerness does not display any change of behavior between the magnetized and the time-crystal phase, it still shows singular behavior at criticality.
Figure 3: Density of nonstabilizerness of the fixed point [$\overline{m_{2}}$, Eq. (18)] and average density of nonstabilizerness along the closed orbits around the fixed point [$\langle\widetilde{m_{2}}\rangle$, Eq. (20)] in the mean-field limit ($\Omega=\omega_{0}/\kappa$). The black dotted line is the numerical fit of Eq. (22). The red dashed line indicates the saturation value, estimated using Eq. (IV).
The SRE density is zero also for $\Omega=0$, where the fixed point is a steady state and is the dark state $|\ket{\psi_{D}}=\ket{\downarrow\downarrow\cdots\downarrow}$ (i. e., a pure state such that $\hat{L}\ket{\psi_{D}}=0$, also a stabilizer state), and $\Omega\to\infty$, where fixed points are the eigenstates of $\hat{H}=\omega_{0}\hat{S}_{x}$. The maximum SRE density is found for $\Omega^{*}_{1}=1/\sqrt{2}$ and $\Omega^{*}_{2}=\sqrt{2}$ and is equal to $\overline{m_{2}}^{*}=2\log 2-\log 3$, which coincides with the stabilizer entropy of the $T$ state |$\ket{T}\sim\ket{\uparrow}+\exp(i\pi/4)\ket{\downarrow}$.
However, it is important to note that, in the BTC phase, the fixed point of the classical MF dynamics is not a steady state, but an oscillation center in phase space [8]. Thus, in this phase it makes more sense to look at the average nonstabilizerness along the closed orbits encircling the fixed point, rather than the fixed point itself.
@@
On the other hand, this is not necessarily true for for the SRE. In this case, the average over the orbit of the limit cycle is not equivalent to the SRE of the fixed point and is generally dependent on the initial condition, which determines the orbit in the semiclassical phase space.
To evaluate $m_{2}$ along these orbits, we proceed numerically. First, we generate $N_{\text{avg}}=1000$ initial conditions of the form $\bm{m}(0)=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, where $\theta\in[0,\pi]$ and $\phi\in[0,2\pi]$ are random variables uniformly distributed in their domains. For each of them, denoted $k$, we solve Eqs. (14) up to a long time $\tau$ and compute the instantaneous nonstabilizerness density $m_{2}^{(k)}(t)$ of this classical trajectory using Eq. (18). In our simulation, we consider $\kappa\tau=1000$. Then, we consider the (long-time) average SRE along the trajectory,
@@
The result of this procedure is shown in Fig. 3 with orange scatter points. Error bars (denoting the standard error, i. e., the standard deviation divided by $\sqrt{N_{\text{avg}}}$) are not visible as they are smaller than the point size. We see that, in the magnetized phase $\Omega<1$, the stabilizer entropy computed by averaging over random initial states coincides with the SRE of the stationary state, whereas for $\Omega>1$, in the BTC phase, the two quantities differ from each other. In either case, the nonstabilizerness density is finite, so the system always has extensive SRE and there are no nonstabilizerness phase transitions associated with the mean-field transition. Yet, the cusp at criticality remains and even becomes more pronounced, as the right derivative of the averaged SRE density at $\Omega_{c}$ becomes larger than the corresponding derivative of the fixed-point stabilizer entropy. Instead, the maximum at $\Omega_{2}^{*}=\sqrt{2}$ is gone, leaving the room to a monotonic growth of $\langle\widetilde{m_{2}}\rangle$ with the driving frequency.
A theoretical argument allows to compute its saturation value. When $\Omega\to\infty$, the oscillation frequency of the magnetization components goes to infinity as well. Thus, the three components will uniformly sample their respective domain and an average over time and initial conditions will be equivalent to an average over the solid angle of the SRE density of the spherical vector $\bm{m}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$. Thus, one can estimate
@@ V Finite-size results
Figure 4: Density of nonstabilizerness of the average state evolving according to the Lindblad equation [Eq. (2)] as function of time, for several system sizes $N$, compared with the mean-field result. (a) $\Omega=\omega_{0}/\kappa=0.5$. (b) $\Omega=1$. (c) $\Omega=2$. The initial condition is the fully polarized state $|\ket{\psi(0)}=\ket{\uparrow\uparrow\dots\uparrow}$.
@@ V.1 Nonstabilizerness of the average state
We now turn our attention to the study of nonstabilizerness in finite-size systems. We start by directly solving the unconditioned Lindblad dynamics of Eq. (2), and we compute $\mathcal{M}_{2}$ of the average state according to Eq. (6). We integrate the Lindblad equation using the 4th order Runge-Kutta method. Due to the strong symmetry of the model, we can represent the entire dynamics in the maximum spin subspace ($S=N/2$), so the density matrix has a linear dimension $D=N+1$ [7].
Our results are shown in Fig. 4, for system sizes up to $N=80$, for the same values of $\Omega$ shown in Fig. 2. In all panels, we see that the density of SRE of the average state depends on the system size. For $\Omega\leq 1$, we see that the long-time density of SRE of the average state coincides with the density of stabilizer entropy calculated using the mean-field approximation (the gray line), whereas for $\Omega>1$ this is not the case. In fact, in this regime the density of SRE decreases with $N$ at long times, in contrast with the finite $m_{2}$ observed in MF dynamics.
@@
Let us consider the left-hand side of Eq. (23). If one takes the thermodynamic limit first, then the dynamics of collective observables is described by Eqs. (14), the system state is always pure, and the behavior of the stabilizer entropy is the one shown in Fig. 2.
@@
The emergence of the time-crystal phase in the weakly dissipative regime is signaled by the fact that the amplitude of the magnetization oscillations decays more slowly as $N$ increases, but they are only persistent for $N\to\infty$: For any finite size, they will eventually decay, yielding a mixed steady state with $\lvert\bm{m}\rvert<1$ and a density of SRE that goes to zero as $N$ grows.
Both these facts are not consistent with the mean-field limit, where $\hat{\rho}$ is a pure state with a finite density of SRE. In particular, see Fig. 5, where we plot the purity of the average state for the same parameters as Fig. 4. Here is evident that, for any finite size, the purity of the average state is smaller than one in the BTC phase, as a consequence of the noncommutativity of the thermodynamic limit and the long-time limit, Eq. (23), and of the existence of residual correlations in the state.
Instead, the two limits in Eq. (23) do commute in the magnetized phase. Here, the purity goes to one, and the steady-state nonstabilizerness converges to its MF limit.
The behavior of the density of SRE of the steady state is shown in Fig. 6(a) as a function of the driving frequency, further confirming that the Lindblad equation does not agree with the MF prediction of a finite density of stabilizer entropy for large $N$. This mismatch is accompanied by a sudden decrease in the purity of the steady state, as shown in Fig. 6(b), confirming the important role of correlations in this region for finite sizes. As long as the steady state is pure ($\Omega\leq 1$), the MF predictions and the nonstabilizerness of the finite-size state evolving through the Lindblad equation [Eq. (2)] agree with each other (compare this with the SRE of the fixed point shown in Fig. 3 for $\Omega\leq 1$). This is no longer the case when $\text{Tr}\,\hat{\rho}^{2}<1$.
Figure 6: (a) Finite-size density of nonstabilizerness of the steady state for different system sizes, in a deterministic Lindblad evolution, as a function of $\Omega=\omega_{0}/\kappa$. (b) Purity of the steady state as a function of $\Omega$.
@@ V.2 Nonstabilizerness of unravelings
Figure 7: Finite-size long-time averaged nonstabilizerness density for different system sizes in different unravelings, see Eq. (5). (a) Quantum Jump ($\mu=0$). (b) $\mu=2$. (b) Quantum State Diffusion ($\mu\to\infty$). We also show the average mean-field SRE $\langle\widetilde{m_{2}}\rangle$, Eq. (20) (orange points).
In this section, we show that the inconsistencies pointed out in the previous section disappear if, instead of computing the nonstabilizerness of the average state, one considers the average stabilizer entropy over ensembles of monitored quantum trajectories.
Indeed, in Fig. 7 we report the average long-time SRE of ensembles of $M=500$ quantum trajectories, in three unravelings. In the three panels, orange scatter points represent the mean-field results discussed in Sec. IV. For $\Omega<1$ they correspond to the steady state, while for $\Omega>1$ to the average over time and random initial conditions. First of all, we see that the SRE density in the $\Omega<1$ phase is essentially unchanged compared to Fig. 6, meaning that each quantum trajectory of the ensemble behaves very closely to the average state in this phase. This is also consistent with the early findings, showing that the variances of collective observables over the ensemble in this phase go to zero for large $N$ [52].
The situation is much different from the average one in the $\Omega>1$ phase, where we observe a finite, size-independent, density of SRE that tells us that the stabilizer entropy of the ensemble trajectories is extensive, as opposed to the nonstabilizerness of the average state $\hat{\rho}$. Moreover, we see that the ensemble average of $m_{2}$, where each trajectory starts from the same initial state but follows a stochastic dynamics due to the random measurements performed by the environment, is essentially identical to the mean-field SRE averaged over many initial states, Eq. (20) (see the orange points in Fig. 7). This observation tells us that the intrinsic randomness of quantum trajectories allows the system to explore its phase space, and that this information is completely hidden at the level of the average state evolving according to the deterministic Lindblad equation.
Moreover, we observe that the nonstabilizerness in all the unravelings is exactly the same, when the system size is sufficiently large, as we have analytically shown in Sec. III.1, in contrast with the behavior of the bipartite entanglement entropy [52]. We see that all the unravelings give the same stabilizer entropy except for finite-size corrections at small sizes, as shown in Fig. 7(b–c) for $\mu=2$ and $\mu\to\infty$.
@@
All these numerical results are consistent with the interpretation discussed in Sec. III.2: for large $N$, the SRE is unraveling-independent because correlations become negligible when computing the expectation values of collective observables. Indeed we can see by comparing the different panels of Fig. 7 that for small $N$ the nonstabilizerness density is not unraveling independent, due to correlations, and that unraveling-independence only kicks in for $N\gtrsim 20$ (empirically).
Figure 9: Histograms showing the probability density of the (a) nonstabilizerness density, (b) entanglement entropy and (c) magnetization in the same ensembles as Fig. 1. Data are binned in 100 bins. Parameters: $N=40$, $\Omega=2$, $\kappa t=8$.
Thus, unraveling the dissipative dynamics in terms of quantum trajectories is able to resolve the inconsistency between finite-size simulations and the thermodynamic limit in the BTC phase, for two reasons. First of all, the states of the ensemble of quantum trajectories are all pure by construction, therefore the SRE is a proper complexity measure for each of them and there are no issues associated with nonpure states like for the stabilizer entropy of the Lindblad steady state in the BTC phase. Second, due to the intrinsic randomness of quantum trajectories, the dynamics of the single trajectories is much different from the dynamics of the Lindblad average state, and averaging over trajectories and time one gets a result well described in the large-size limit by the average over time and random initial states of MF trajectories.
@@
The same considerations apply to the density of SRE as well, as shown in Fig. 8(b) for the same parameters. Also in this case, the SRE dynamics along single trajectories more closely resembles the dynamics in the MF limit, compared to the Lindblad nonstabilizerness where instead oscillations decay over time and the stabilizer entropy goes to zero. Moreover, since the SRE is a nonlinear function, its ensemble average over quantum trajectories does not coincide with the SRE of the average state. Therefore we think it is reasonable that its value, long-time averaged and also averaged over the pure states of the ensemble, would resemble the MF limit (averaged over many initial states) more than the nonstabilizerness of the average state for finite $N$.